OpenStudy (anonymous):

find the average value of the function f(x)=1+x^2 on interval [-1,2] i need help on last step where you plug the -1, 2 in the soultion

OpenStudy (anonymous):

you may try it now.

OpenStudy (anonymous):

write down the formula first.

OpenStudy (anonymous):

i know but i could not know how mutiply...1/3[x+x^3/3] to the interval -1, 2

OpenStudy (anonymous):

you are talking about the average value of the function, use this formula : [f(x+h)-f(x)]/h , wher h=b-a

OpenStudy (anonymous):

i think more like 1/b-1 intergral a to b f(x) dx

OpenStudy (anonymous):

you are talking then about the mean value theorem by integration

OpenStudy (anonymous):

no the average value function...just do the steps for me n illl figure out where im stuck

OpenStudy (anonymous):

please

OpenStudy (anonymous):

Average value of any function in the given interval is the integral of that function in that given interval divided by the length of that interval.

OpenStudy (anonymous):

h=2-(-1) =3 ===> f(x+3) = 1+((x+3)^2 =1+x^2+6x+9 , f(x)=1+x^2

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

it depends on the question if it's in the integation or in the deffrentiation

OpenStudy (anonymous):

but both way work

OpenStudy (anonymous):

intrgation

OpenStudy (anonymous):

so just find the definite integration / 3

OpenStudy (anonymous):

x^3/3

OpenStudy (anonymous):

average value of the function f(x)=1+x^2 = (x+(x^3)/3)/(2-(-1)) with limits -1 to 2 = 2

OpenStudy (anonymous):

[x^3/3] (from -1 to 2) = 8/3 -(-1/3) =9/3=3

OpenStudy (anonymous):

thank you nagavishnu now i undersatnd thanks thats where i got stuck on

OpenStudy (anonymous):

how you mulitply it

OpenStudy (anonymous):

I think you missed the integration itself , it should as nagavi posted

OpenStudy (anonymous):

\[f_{avg}={1 \over b-a}\int\limits_{a}^{b}f(x)dx={1 \over 2-(-1)}\int\limits_{-1}^{2}(1+x^2)dx={1 \over 3}(x+{x^3 \over 3})_{-1}^{2}\] \[={1 \over 3}(2+{8 \over 3}+1+{1 \over 3})=4\] I hope I din't make any mistakes.

OpenStudy (anonymous):

you did it as x^3/3 , whearas it is (x+(x^3)/3)

OpenStudy (anonymous):

the answer is 2 not 4

OpenStudy (anonymous):

it's done now

OpenStudy (anonymous):

yes , the final answer = 6/3=2

OpenStudy (anonymous):

Just a typo.. It's \(\frac{1}{3}(2+\frac{8}{3}+1+\frac{1}{3})=\frac{6}{3}=2\).